OpenStudy (anonymous):

area of triangle flag. as part an out door education course. trisha needs to make a bright-colored triangular flag with an area of at least 3 ft^2 what heights can the triangle be if base 1 1/2

OpenStudy (anonymous):



OpenStudy (unklerhaukus):

A = 1/2 x B x H A=3, B= (3/2) H=2A / B =2x(3) /(3/2) = 4

OpenStudy (unklerhaukus):

or anything greater